Pablo owns one share of stock of Blue Eagle Media and one share of stock of Violet Sky Consulting. The total value of his holdings is 254.72 dollars. Both stocks pay annual dividends that are expected to continue forever. The expected return for Blue Eagle Media stock is 15.57 percent and its annual dividend is expected to remain at 9.38 dollars forever. What is the expected return for Violet Sky Consulting stock if its next dividend is expected to be 9.51 dollars and all subsequent dividends are expected to grow by 5.17 percent annually? The next dividend for both firms’ stocks will be paid in one year. Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

Value of blue eagle media = Annual dividend / required rate

Value of blue eagle media = 9.38 / 0.1557

Value of blue eagle media = $60.2441

Value of violet sky = 254.72 - 60.2441 = 194.4759

Expected return of violet sky = (D1 / share price) + growth rate

Expected return of violet sky = (9.51 / 194.4759) + 0.0517

Expected return of violet sky = 0.1006